Can you please show me the steps to my homework problem? Its y=x+8
-x+3y=10
k thanks.

Answer by nerdybill(7384)

We're going to apply the "elimination" method:
y=x+8
-x+3y=10
.
Move all the x's and y's to the left:
-x+ y= 8
-x+3y=10
.
Multiply both sides of the TOP equation by -1. Then combine BOTH equations together to eliminate the x term:
x- y=-8
-x+3y=10
-----------
2y=2
y = 1
.
Now, we substitute y=1 into:
y=x+8
1=x+8
-7=x
.
solution: (x,y) = (-7,1)